Operations Report – January 25th

SOL: 4
Name of person filing report: Scott Dorrington
Non-nominal systems: None
Notes on non-nominal systems: None
Spirit rover used:
Hours: 242.6
Beginning charge: 100%
Ending charge: 56%
Currently charging: No
Opportunity rover used:
Hours: 156.8
Beginning charge: 100%
Ending charge: 58%
Currently charging: No
Curiosity rover used:
Hours: 264.5
Beginning charge: 100%
Ending charge: 46%
Currently charging: No
Perseverance rover used:
Hours: 282.4
Beginning charge: 99%
Ending charge: 38%
Currently charging: No
General notes on rovers: Mirrors removed previous night.
Summary of Hab operations:
WATER USE: 13.1 gallons
Water (static tank): 380.1 gallons (depth 13 inches)
Static tank pipe heater (on or off): on
Static tank heater (On or off): on
Toilet tank emptied: yes (indicator light has not changed from 2/3 full – visually monitoring)
Summary of internet: Nominal
Summary of suits and radios: Suit #8 was not used today, as it was noted to have two screws loose at the charging port. These came out when one of the crew attempted to remove the power plug. (See pictures before and after repair). I screwed them both back in, but found the top screw would not tighten, just keeps turning. The plug is secured, and can function for plugging and unplugging, but may need extra care and monitoring when plugging/unplugging. We are prepared to stand suit #8 down for now until further advice from mission support.
Summary of GreenHab operations: put notes here
WATER USE: 5 gallons
Heater: Space heater on overnight
Supplemental light: light on all day
Harvest: None
Summary of Science Dome operations: Crew visited Science Dome in search of vinegar and hydrochloric acid for use in rock sampling in the field. Identified chemicals, but not used (could be part of ongoing experiments).
Dual split: Nominal
Summary of RAM operations: Not used
Summary of any observatory issues: None
Summary of health and safety issues: Medical forehead thermometer was tested but shown to be not accurate to expected body temperatures – out by 4 to 5 C.
Questions, concerns and requests to Mission Support:
Crew were interested in the vinegar or hydrochloric acid (10% HCl) in the Science Dome for rock testing of returned samples. Inquiring if this is from an ongoing experiment, or if we are allowed to use it. Acknowledge that handbook requires crew to bring their own chemicals for testing, so ok if not allowable.

Copyright © The Mars Society. All rights reserved. | Main Site